A TELIVISION HAS A 20 INCH DIAGONAL SCREEN AND A 12 INCH HEIGHT . WHAT IS THE LENGH OF THE TELEVISION?
Answer by swincher4391(1107)

Pythagorean Theorem
x^2 + 12^2 = 20^2
x^2 + 144 = 400
x^2 = 400 -144
x^2 = 256
x = 16 inches